OpenStudy (anonymous):

why is cos°=(|180-°)?

OpenStudy (lgbasallote):

why is what?

OpenStudy (lgbasallote):

is there something missing there?

OpenStudy (anonymous):

why is cos°=(|180-°)

OpenStudy (anonymous):

\[\cos \theta=(|180-\theta)\]